# CausalDynamics: A large-scale benchmark for structural discovery of dynamical causal models

### v1

@@ D.1 Summary
The final distribution of our preprocessed dataset is illustrated in Figure 13. For the SDE cases, we combine all dynamics with noise level $\delta>0$. The nonlinear and periodic attributes refer to graphs where at least one of the root nodes represents only a simple dynamics or also mixed with periodic forcing.
@@ D.2 Hierarchy of Differential Equations
Coupled. We then couple the simple systems and perform the following combination of experiments: (i) varying the Langevin noise amplitude ($\delta\in\{0.0,0.5,1.0,1.5,2.0\}$, (ii) mimicking unobserved confounding scenarios, (iii) changing the number of coupled nodes ($N_{node}\in\{3,5,8\}$), (iv) introducing periodic drivers (Coupled: Periodic) at equal ratio as dynamical system drivers or only having the latter, i.e., nonlinear chaotic drivers (Coupled: Nonlinear), (v) time-signal standardization, (vi) time-lags ($\tau\in\{0,1\}$). Though 4800 coupled graphs are supposed to be generated, we keep 4745 of them, as the rest ($<1\%$) exhibit undesired outcomes, e.g., trajectories diverge, even after multiple retries. Finally, the success rate for evaluating coupled cases is $99\%$ (or 33208 out of 33215 possible evaluations). Similar to the simple case, one of the most common failure reasons involves the singular value of the true adjacency matrix where score metrics such as AUROC fail.
@@ E.1 Baselines
As part of our benchmark, we deploy 7 methods. We also summarize the final set of hyperparameters after a rudimentary search and attempt to set them to their default configurations. Nevertheless, we find that methods with little to no fine-tuning often outperform those that do so intensively. In all algorithms, we set the maximum lag $\tau_{\text{max}}=1$, unless otherwise specified. This choice is motivated by the fact that dynamical systems governed by differential equations typically evolve based on a single discretization step, which suffices to capture the system’s causal structure in the resulting time series.
PCMCI+ (Peter Clark Momentary Conditional Independence) [40] extends the PC-algorithm to multivariate time series by first preselecting candidate parents and then applying momentary conditional independence (MCI) tests across all lags to estimate contemporaneous and lagged edges under controlled false discovery. We set the maximum lagged time ($\tau_{max}=1$) and the critical p-value ($\alpha_{crit}=0.01$).
F-PCMCI (Filtered PCMCI) [69] extends PCMCI+ by incorporating transfer-entropy-based feature selection to prefilter candidate parents before momentary conditional independence testing, improving scalability and robustness in high-dimensional time series. We set the maximum lagged time ($\tau_{max}=1$) and the critical p-value ($\alpha_{crit}=0.01$).
@@
DYNOTEARS [20] adopts a score-based structural VAR (SVAR) formulation to jointly estimate contemporaneous and time-lagged weight matrices by minimizing a penalized least-squares objective expended with a differentiable penalty that forbids any directed cycles in the instantaneous effects. We set the maximum lagged time to $\tau_{max}=1$.
@@
TSCI (Tangent Space Causal Inference) [73] builds on the idea that nonlinear dynamics locally resemble linear systems by estimating the data manifold’s tangent spaces. TSCI first learns a continuous vector-field model for each variable’s dynamics, e.g., via neural ODEs or Gaussian processes, then applies Convergent Cross Mapping (CCM) within each tangent to test for causal influence by assessing how well the state-space of one variable predicts another. This yields a single, interpretable causal graph for deterministic dynamical systems. We set the maximum lagged time ($\tau_{max}=1$), the delay embedding dimension ($h_{embed}=2$), and the correlation threshold ($\rho_{crit}=0.8$).
@@
Neural GC (Neural Granger Causality) [5] employs sparse-input multilayer perceptrons (cMLP) and long short-term memory (cLSTM) to model nonlinear autoregressive dependencies in multivariate time series, using input-weight regularization to identify directed Granger-causal links. In our evaluation, we focus solely on the cLSTM variant. We set the maximum lagged time ($\tau_{max}=1$), number of hidden LSTM dimension ($h_{lstm}=16$), with a learning rate of ($lr=10^{-3}$) fitted over 100 epochs.
@@
CUTS+ [68] builds on the Granger-causality framework of CUTS [57] by combining a two-stage coarse-to-fine strategy with a message-passing graph neural network to discover causal structure in high-dimensional, irregularly-sampled time series. First, lightweight Granger-causality tests pre-filter candidate parents for each variable, dramatically reducing the search space. Then, a GNN simultaneously imputes missing or unevenly spaced data and learns a sparse adjacency matrix via a penalized reconstruction loss that incorporates temporal encoding. This alternating imputation–learning loop makes CUTS+ both scalable and robust to missing data. We set the maximum lagged time ($\tau_{max}=1$), the number of hidden MLP dimension ($h_{mlp}=16$), the number of gated recurrent unit (GRU) ($n_{gru}=1$), with a learning rate of $lr=10^{-3}$ fitted over 10 epochs.
@@ E.2 Metrics
We describe the metrics used in this work, including AUROC [88, 89] and AUPRC [90].
@@
[…]
which defines the probability that a true edge ranks above a false one, i.e., $0.5$ means random, $<0.5$ means worse than random, $1.0$ means a perfect prediction.
@@
[…]
which unlike AUROC depends on the edge-density $\frac{\mathcal{E}}{n(n-1)}$ and thus more accurately reflects performance in the sparse-graph regime. For AUROC the optimal score is 1, i.e., perfect prediction.

### v2

@@ D.1 Summary
The final distribution of our preprocessed dataset is illustrated in Figure 13. For the SDE cases, we combine all dynamics with noise level $\delta>0$. The nonlinear and periodic attributes refer to graphs where at least one of the root nodes represents only a simple dynamics or also mixed with periodic, nonlinear forcing.
@@ D.2 Hierarchy of Differential Equations
Coupled. We then couple the simple systems and perform the following combination of experiments: (i) varying the Langevin noise amplitude ($\delta\in\{0.0,0.5,1.0,1.5,2.0\}$, (ii) mimicking unobserved confounding scenarios, (iii) changing the number of coupled nodes ($N_{node}\in\{3,5,10\}$), (iv) introducing periodic drivers (Coupled: Periodic) at equal ratio as dynamical system drivers or only having the latter, i.e., nonlinear chaotic drivers (Coupled: Nonlinear), (v) time-signal standardization, (vi) time-lags ($\tau\in\{0,1\}$), (vii) nonlinear edge-level activations. Though 18000 coupled graphs are supposed to be generated, we keep 14096 of them, as the rest exhibit undesired outcomes, e.g., trajectories diverge, even after multiple retries.
@@ E.1 Baselines
As part of our benchmark, we deploy 10 methods. We also summarize the final set of hyperparameters after a comprehensive search on a held-out dataset from the simple case, and attempt to set them to their default configurations. Nevertheless, we find that methods with little to no fine-tuning often outperform those that do so intensively. In all algorithms, we set the maximum lag $\tau_{\text{max}}=1$, unless otherwise specified. This choice is motivated by the fact that dynamical systems governed by differential equations typically evolve based on a single discretization step, which suffices to capture the system’s causal structure in the resulting time series.
PCMCI+ (Peter Clark Momentary Conditional Independence) [44] extends the PC‐algorithm to multivariate time series by first preselecting candidate parents and then applying momentary conditional independence (MCI) tests across all lags to estimate contemporaneous and lagged edges under controlled false discovery. We set the maximum lagged time ($\tau_{max}=1$) and the critical p-value ($\alpha_{crit}=0.01$).
F-PCMCI (Filtered PCMCI) [73] extends PCMCI+ by incorporating transfer‐entropy‐based feature selection to prefilter candidate parents before momentary conditional independence testing, improving scalability and robustness in high‐dimensional time series. We set the maximum lagged time ($\tau_{max}=1$) and the critical p-value ($\alpha_{crit}=0.01$).
@@
DYNOTEARS [22] adopts a score‐based structural VAR (SVAR) formulation to jointly estimate contemporaneous and time‐lagged weight matrices by minimizing a penalized least‐squares objective expended with a differentiable penalty that forbids any directed cycles in the instantaneous effects. We set the maximum lagged time to $\tau_{max}=1$.
@@
TSCI (Tangent Space Causal Inference) [79] builds on the idea that nonlinear dynamics locally resemble linear systems by estimating the data manifold’s tangent spaces. TSCI first learns a continuous vector‐field model for each variable’s dynamics, e.g., via neural ODEs or Gaussian processes, then applies Convergent Cross Mapping (CCM) within each tangent to test for causal influence by assessing how well the state‐space of one variable predicts another. This yields a single, interpretable causal graph for deterministic dynamical systems. We set the maximum lagged time ($\tau_{max}=1$), the delay embedding dimension ($h_{embed}=2$), and the correlation threshold ($\rho_{crit}=0.8$).
@@
Neural GC (Neural Granger Causality) [5] employs sparse‐input multilayer perceptrons (cMLP) and long short-term memory (cLSTM) to model nonlinear autoregressive dependencies in multivariate time series, using input‐weight regularization to identify directed Granger‐causal links. In our evaluation, we focus solely on the cLSTM variant. We set the maximum lagged time ($\tau_{max}=1$), number of hidden LSTM dimension ($h_{lstm}=16$), with a learning rate of ($lr=10^{-3}$) fitted over 100 epochs.
@@
CUTS+ [72] builds on the Granger‐causality framework of CUTS [61] by combining a two‐stage coarse-to-fine strategy with a message-passing graph neural network to discover causal structure in high-dimensional, irregularly-sampled time series. First, lightweight Granger-causality tests pre-filter candidate parents for each variable, dramatically reducing the search space. Then, a GNN simultaneously imputes missing or unevenly spaced data and learns a sparse adjacency matrix via a penalized reconstruction loss that incorporates temporal encoding. This alternating imputation–learning loop makes CUTS+ both scalable and robust to missing data. We set the maximum lagged time ($\tau_{max}=1$), the number of hidden MLP dimension ($h_{mlp}=16$), the number of gated recurrent unit (GRU) ($n_{gru}=1$), with a learning rate of $lr=10^{-3}$ fitted over 10 epochs.
RCD (Repetitive Causal Discovery) [77] extends LiNGAM to settings with latent confounders under the linear, non-Gaussian, acyclic model (LiNGAM) assumption. Classic LiNGAM [75] identifies a unique DAG by exploiting non-Gaussianity and independence of error terms, but its basic form assumes causal sufficiency. RCD relaxes this by repeatedly partitioning variables into (approximate) causal orders and pruning edges using statistical tests, while allowing for latent common causes. Concretely, RCD alternates between: (i) screening and ordering steps that use correlation/independence diagnostics to propose parent sets under linear models with non-Gaussian residuals, and (ii) refinement steps that remove spurious links via independence checks on residuals (testing that, in the correct direction, residuals are independent of putative causes) and normality checks to ensure non-Gaussianity (a key LiNGAM identifiability condition). The repetitive loop iteratively updates ordering and adjacency until convergence, yielding a sparse adjacency estimate that is robust to latent confounding under the RCD model assumptions. We set the maximum lagged time ($\tau_{max}=1$) and the critical p-value ($\alpha_{crit}=0.01$).
GRaSP (Greedy Sparsest Permutations) [74] is a permutation-based causal discovery method that learns DAGs by searching over variable orderings. The method is grounded in the permutation-DAG correspondence: any variable ordering defines a unique minimal DAG consistent with the observed conditional independencies. Identifying the true DAG is then equivalent to finding the sparsest permutation, i.e., the one that yields the fewest edges. GRaSP employs a greedy local search strategy to approximate this combinatorial optimization. Starting from a random initial ordering, it iteratively applies adjacent swaps of variables. At each step, the algorithm updates the candidate DAG using conditional independence tests and accepts swaps that reduce the number of edges, gradually moving toward a sparser representation. This approach avoids the exhaustive search over all $n!$ permutations, making it scalable to moderately high-dimensional problems. By focusing on sparsity, GRaSP is particularly effective when the true underlying causal graph is relatively sparse and conditional independencies can be reliably tested. We set the maximum lagged time ($\tau_{max}=1$).
TCDF (Temporal Causal Discovery Framework) [68] uses convolutional neural networks (CNN) with an attention mechanism to detect lagged causal relationships in multivariate time series, combining temporal filters and statistical pruning to handle nonlinear, high-dimensional, and noisy data. We set the maximum lagged time ($\tau_{max}=1$), the CNN kernel size ($h_{cnn}=4$), the dilation coefficient of 4 with a learning rate of $lr=10^{-2}$ fitted over 100 epochs.
@@ E.2 Metrics
We describe the metrics used in this work, including AUROC [94, 95], AUPRC [96], and SHD [97, 98].
@@
$\text{AUROC}=\int_{0}^{1}\text{TPR}\bigl(\text{FPR}^{-1}(u)\bigr)\,du,$ (19)
which defines the probability that a true edge ranks above a false one, i.e., $0.5$ means random, <0.5<0.5 means worse than random, $1.0$ means a perfect prediction.
@@
$\text{AUPRC}=\int_{0}^{1}\mathrm{Precision}\bigl(\text{Recall}^{-1}(r)\bigr)\,dr.$ (21)
which unlike AUROC depends on the edge‐density $\frac{\mathcal{E}}{n(n-1)}$ and thus more accurately reflects performance in the sparse‐graph regime. For AUROC the optimal score is 1, i.e., perfect prediction.
Finally, SHD is defined as the minimum number of edge additions, deletions, or reversals required to transform $\hat{\mathcal{G}}$ into $\mathcal{G}$:
$\mathrm{SHD}(\hat{\mathcal{G}},\mathcal{G})=\bigl|\{(i,j)\in\mathcal{E}\setminus\hat{\mathcal{E}}\}\bigr|+\bigl|\{(i,j)\in\hat{\mathcal{E}}\setminus\mathcal{E}\}\bigr|+\#\{(i,j):(i,j)\in\hat{\mathcal{E}},(j,i)\in\mathcal{E}\}.$ (22)
which a smaller SHD indicates a closer match to the ground truth, with 0 corresponding to exact recovery. Unlike AUROC and AUPRC, which are ranking-based metrics, SHD directly evaluates the correctness of the learned graph structure.
Appendix F Additional results
F.1 Time discretisation ablation
We investigate the sensitivity of causal discovery algorithms to different sampling frequencies by sub-sampling every 1st, 5th, or 10th time step. As an illustration, we consider one of the most difficult experiments: coupling=nonlinear_noise=2.00_systems=10_confounder=True_standardize=True_timelag=1. Results are shown in Table 5. Overall, we find that reducing sampling frequency tends to improve detection accuracy. This is potentially due to the reduction of data redundancy that can noise the algorithm.
Table 5: Ablating time discretisation: AUROC, AUPRC, and SHD scores at different sampling frequencies (every 1st, 5th, or 10th step).
Model AUROC (1 / 5 / 10) AUPRC (1 / 5 / 10) SHD (1 / 5 / 10)
PCMCI+ 0.486 / 0.550 / 0.553 0.382 / 0.415 / 0.425 51.9 / 44.4 / 40.5
F-PCMCI 0.547 / 0.557 / 0.566 0.406 / 0.418 / 0.429 45.7 / 44.2 / 38.9
VARLiNGAM 0.519 / 0.518 / 0.548 0.389 / 0.389 / 0.407 51.9 / 51.3 / 47.2
DYNOTEARS 0.546 / 0.539 / 0.574 0.426 / 0.416 / 0.444 38.9 / 41.6 / 37.9
NGC-LSTM 0.497 / 0.500 / 0.500 0.377 / 0.378 / 0.378 56.0 / 34.0 / 34.0
TSCI 0.533 / 0.576 / 0.567 0.416 / 0.458 / 0.443 41.4 / 38.6 / 36.4
CUTS+ 0.495 / 0.500 / 0.500 0.376 / 0.378 / 0.378 39.0 / 34.0 / 34.0
RCD 0.500 / 0.501 / 0.505 0.378 / 0.382 / 0.385 34.0 / 34.1 / 34.1
GRaSP 0.498 / 0.484 / 0.490 0.385 / 0.379 / 0.384 35.8 / 36.2 / 36.2
TCDF 0.500 / 0.500 / 0.500 0.378 / 0.378 / 0.378 34.0 / 34.0 / 34.0
F.2 Partial observability ablation
We also test algorithm performance under partial versus full observability, by sampling only one variable per multidimensional node versus averaging across all variables at each node. As an illustration, we consider one of the most difficult experiments: coupling=nonlinear_noise=2.00_systems=10_confounder=True_standardize=True_timelag=1. Results are shown in Table 6. We find that fully observing the system yields a modest improvement in detection performance. One possible explanation is that, even when some variables are hidden, their influence still propagates through the variables we do observe, so the loss of information is less severe than it might appear.
Table 6: Ablating partial observability: AUROC, AUPRC, and SHD scores for partially vs. fully observed systems.
Model AUROC (partial / full) AUPRC (partial / full) SHD (partial / full)
PCMCI+ 0.486 / 0.500 0.382 / 0.386 51.9 / 50.2
F-PCMCI 0.545 / 0.546 0.405 / 0.406 45.8 / 47.4
VARLiNGAM 0.519 / 0.503 0.389 / 0.381 51.9 / 51.7
DYNOTEARS 0.546 / 0.559 0.426 / 0.426 38.9 / 37.2
NGC-LSTM 0.499 / 0.500 0.377 / 0.378 55.9 / 34.0
TSCI 0.533 / 0.552 0.416 / 0.421 41.4 / 38.5
CUTS+ 0.496 / 0.500 0.376 / 0.378 39.2 / 34.0
RCD 0.500 / 0.499 0.378 / 0.377 34.0 / 34.0
GRaSP 0.498 / 0.499 0.385 / 0.388 35.8 / 36.4
TCDF 0.500 / 0.500 0.378 / 0.378 34.0 / 34.0
F.3 Edge-level activation ablation
We also perform additional experiment where the edge-level activation functions vary. For each identical experiment_id, we pass a new argument activation=mixed, which by default randomly samples activations_names=[‘‘identity’’, ‘‘sin’’, ‘‘sigmoid’’, ‘‘tanh’’, ‘‘relu’’] as the graph’s edge-level activation. We compare results for the following experiments with coupling=nonlinear_noise=0.00_systems=3_confounder=False_standardize=False_timelag=0:
  1. 1.
    Linear activation (identity)
  2. 2.
    Mixed activation
Table 7: Ablating different (linear vs. mixed nonlinear) activations: AUROC, AUPRC, and SHD scores.
Model AUROC (linear / mixed) AUPRC (linear / mixed) SHD (linear / mixed)
CUTS+ 0.500 / 0.478 0.378 / 0.417 37.200 / 45.667
DYNOTEARS 0.532 / 0.646 0.407 / 0.541 39.500 / 30.167
F-PCMCI 0.546 / 0.604 0.405 / 0.493 46.650 / 36.333
GRaSP 0.521 / 0.488 0.396 / 0.437 35.500 / 42.667
NGC-LSTM 0.498 / 0.503 0.377 / 0.428 55.900 / 51.278
PCMCI+ 0.526 / 0.560 0.398 / 0.462 47.900 / 41.167
RCD 0.500 / 0.495 0.378 / 0.424 34.000 / 38.333
TCDF 0.500 / 0.491 0.378 / 0.430 34.000 / 41.167
TSCI 0.537 / 0.689 0.414 / 0.601 39.850 / 24.889
VARLiNGAM 0.503 / 0.554 0.380 / 0.458 51.700 / 43.667
We find that mixed activations tend to improve causal discovery performance, potentially because directionality is more detectable in nonlinear cases, i.e., the signal is less invertible (see LiNGAM paper [75]). This effect is especially pronounced for deep learning–based methods such as TSCI, which shows an improvement of the SHD score by 40% (Table 7). This suggests that DL-based methods are better at capturing nonlinearities in the data.